OpenStudy (anonymous):

Convert the rectangular equation to polar form. 2x-3y=5

zepdrix (zepdrix):

\[\Large\bf\sf x=r \cos \theta\]\[\Large\bf\sf y=r \sin \theta\]Plug em' in.

zepdrix (zepdrix):

We usually write a polar equation in the form r=

zepdrix (zepdrix):

Try to solve for r if you can.

OpenStudy (anonymous):

final answer i got: 5/ (2cosΘ - 3sinΘ) is that right

zepdrix (zepdrix):

Mmm yes good job :)